# On codes induced from Hadamard matrices 111 Keywords: Hadamard, code, linear, convolutional, self-dual, dual-containing, quantum code, linear complementary AMS Classification 2020: 94B05, 94B10, 15B99

### v1

@@ 1 Introduction
$C^{\perp}$ denotes the dual of the code $\mathcal{C}$. See section 1.1 for precise definition of dual of a convolutional code. $\mathcal{C}$ is a dual-containing (DC) code if $\mathcal{C}\cap C^{\perp}=\mathcal{C}^{\perp}$; $\mathcal{C}$ is a linear complementary dual (LCD) code if $\mathcal{C}\cap C^{\perp}=0$. $\mathcal{C}$ is a self-dual code provided $\mathcal{C}^{\perp}=\mathcal{C}$ and is an important type of dual-containing code.
Dual-containing codes, which include self-dual codes, can be used to construct quantum error-correcting codes (QECC) by the CSS method [7, 9, 8]; here convolutional quantum error-correcting codes are constructed in this way from Hadamard matrices.
A Hadamard matrix is an $n\times n$ matrix $H$ with entries ±$\pm 1$ satisfying $HH\text{${}^{\text{T}}$}=nI_{n}$. Such a matrix can only exist for $n=2$ or $n=4m$ for a positive integer $m$, [5] Theorem 18.1; see [6] for a beautifully written book on Hadamard matrices. Here these ±$\pm 1$ entries may be considered as elements in a general field.
The Walsh-Hadamard codes can be formed from Walsh-Hadamard matrices of size $2^{k}\times 2^{k}$ by equating the entries $(-1)$ to 00 and then forming binary codes from the $(k+1)$ linearly independent rows remaining. Using the unit-derived methods on a general Hadamard $n\times n$ matrix gives much more scope, arbitrary rates, good distances, required types but also both linear block and convolutional codes are formed. Several algorithms exist for decoding convolutional codes, the most common ones being the Viterbi algorithm and the sequential decoding algorithm.
@@
Applications are obtained by applying the Propositions and Algorithms to particular cases. Once the Hadamard matrix is formed, the Propositions and Algorithms may then be applied. For large lengths a problem may be to determine the distance but this is not necessary for applications.
@@
    self-dual $[20,10,8]_{5}$ codes;
@@
    $[40,20,12]_{3}$ self-dual codes given directly in systematic form, see Proposition 2.2.
@@
  • LCD $[28,16,6]_{3},[28,12,9]_{5}$ codes;
  • Convolutional self-dual $(28,14,14;1,12)_{3^{2}}$ codes over $GF(3^{2})$;
  • DC convolutional $(28,18,10;1,8)_{3^{2}}$ codes over $GF(3^{2})$;
  • quantum codes of length $28$, distance $8$, rate $\frac{8}{28}$ over $GF(3^{2})$;
  • self-dual convolutional $(28,14,14;1,16)_{5}$ codes over $GF(5)$
  • DC convolutional $(28,16,12;1,14)_{5}$ codes over $GF(5)$;
  • $[28,14,9]_{7}$ self-dual codes.
Generally from $H(n)$ with $p\nmid n$, self-dual $(n,\frac{n}{2},\frac{n}{2};1,d)$ convolutional codes and DC $(n,r,n-r;1;d)$, $r>\frac{n}{2}$, convolutional codes are formed. In prototype example 2.9, it is shown how the different types of LCD, self-dual, DC, quantum, linear block and convolutional codes may be derived for a small Hadamard matrix case. Self-dual codes over $GF(p)$ can often be obtained from a Hadamard matrix of size $n$ when $p\mid n,p\neq 2$, see Proposition 2.8. For example self-dual $[12k,6k,d]_{3}$ codes are produced from Paley-Hadamard matrices of size $12k$, see section 2.1.
@@ 1.1 Additional notation and background
The notation for linear block codes is standard and may be found in [1, 2, 3, 4] and many others. $GF(q)$ denotes the finite field with $q$ elements and $\mathbb{Z}_{m}$ denotes the integers modulo $m$; in particular $\mathbb{Z}_{p}=GF(p)$ for a prime $p$. A $[n,r,d]$ code denotes a linear block code of length $n$, dimension $r$, and (minimum) distance $d$; the rate is $\frac{r}{n}$. A $[n,r,d]_{q}$ code denotes a linear block code of length $n$, dimension $r$ and distance $d$ over the field $GF(q)$.
Different equivalent definitions for convolutional codes are given in the literature. The notation and definitions used here follow that given in [21, 40, 22]. A rate $\frac{k}{n}$ convolutional code with parameters $(n,k,\delta)$ over a field $\mathcal{F}$ is a submodule of $\mathcal{F}[z]^{n}$ generated by a reduced basic matrix $G[z]=(g_{ij})\in\mathcal{F}[z]^{r\times n}$ of rank $r$ where $n$ is the length, $\delta=\sum_{i=1}^{r}\delta_{i}$ is the degree with $\delta_{i}=\max_{1\leq j\leq r}{\deg g_{ij}}$. Then $\mu=\max_{1\leq i\leq r}{\delta_{i}}$ is known as the memory of the code and the code is then given with parameters $(n,k,\delta;\mu)$. The parameters $(n,r,\delta;\mu,d_{f})$ are used for such a code with free (minimum) distance $d_{f}$. Further $(n,r,\delta;\mu,d_{f})_{q}$ is used to specify that the code is over the field $GF(q)$.
@@
Let $G(z)$ be the generator matrix for a convolutional code $\mathcal{C}$ with memory $m$. Suppose $G(z)H\text{${}^{\text{T}}$}(z)=0$, so that $H\text{${}^{\text{T}}$}(z)$ is a control matrix, and then $H(z^{-1})z^{m}$ generates the convolutional dual code of $\mathcal{C}$, see [23] and [28]. This is also known as the module-theoretic dual code. [^3] In convolutional coding theory, the idea of dual code has two meanings. The other dual convolutional code defined is called the sequence space dual; the generator matrices for these two types are related by a specific formula. The code is then dual-containing provided the code generated by $H(z^{-1})z^{m}$ is contained in the code generated by $G(z)$.
@@ 2 Linear block and convolutional codes induced from Hadamard matrices
The Walsh-Hadamard binary linear block codes $[2^{k},k,2^{k-1}]_{2}$ and $[2^{k},k+1,2^{k-1}]_{2}$ have very small rate but have found use probably on account of the distances and decoding methods available.
Codes from general Hadamard matrices as now described allow much more scope with much better rates, good distances, required types and both linear block and convolutional codes may be formed.
@@
Let $H$ be a Hadamard matrix of size $n$ so that $HH\text{${}^{\text{T}}$}=nI_{n}$ and $n\neq 0$ in a field $\mathcal{F}$. Suppose $H$ has the form $H=\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)$ implying $\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)\left(\begin{smallmatrix}A\text{${}^{\text{T}}$}&B\text{${}^{\text{T}}$}\end{smallmatrix}\right)=nI_{n}$ where $A$ has size $r\times n$ and $B$ has size $(n-r)\times n$ with $r>(n-r)$. Let $t=(2r-n)$ and define $B_{1}=\bigl{(}\begin{smallmatrix}0_{t\times n}\\ B\end{smallmatrix}\bigr{)}$. Then $G(z)=A+iB_{1}z$ generates a convolutional dual-containing $(n,r,n-r;1,d)$ code $\mathcal{C}$ where $i=\sqrt{-1}$ in $\mathcal{F}$ or in a quadratic extension of $\mathcal{F}$.
@@
Define $0_{t}=0_{t\times n}$. Thus $B_{1}=\left(\begin{smallmatrix}0_{t}\\ B\end{smallmatrix}\right)$ is an $r\times n$ matrix. Now $A\text{${}^{\text{T}}$}$ is an $n\times r$ matrix and thus has the form $A\text{${}^{\text{T}}$}=(X,C_{1})$ where $C_{1}$ has size $n\times(n-r)$ and $X$ has size $n\times(2r-n)$. As $AA\text{${}^{\text{T}}$}=nI_{r}$ then $AC_{1}=n\begin{pmatrix}0_{(2r-n)\times(n-r)}\\ I_{(n-r)\times(n-r)}\end{pmatrix}$ and also $B_{1}B=n\begin{pmatrix}0_{(2r-n)\times(n-r)}\\ I_{(n-r)\times(n-r)}\end{pmatrix}$. Now $A=\left(\begin{smallmatrix}X\text{${}^{\text{T}}$}\\ C_{1}\text{${}^{\text{T}}$}\end{smallmatrix}\right),B_{1}=\left(\begin{smallmatrix}0_{t}\\ D\text{${}^{\text{T}}$}\end{smallmatrix}\right)$. Then $(A+iB_{1}z)(iB\text{${}^{\text{T}}$}+C_{1}z)=0$ so $H\text{${}^{\text{T}}$}(z)=(iB\text{${}^{\text{T}}$}+C_{1}z)$ is a control matrix and $H(z^{-1})z=C_{1}+iB$ generates the dual of the code. The code generated by $C_{1}+iB$ is easily seen to be contained in the code generated by $(A+iB_{1})$ and so the code generated by $(A+iB_{1})$ is dual-containing. That the code is non-catastrophic follows in a similar manner to the proof in Proposition 2.3.
@@
Then $G(z)=\left(\begin{smallmatrix}A&B\end{smallmatrix}\right)+i\left(\begin{smallmatrix}\underline{0}\\ C\end{smallmatrix}\right)$ defines a dual-containing convolutional $(n,r,n-r;1,d)$ code where $i=\sqrt{-1}$ in $\mathcal{F}$ or in a quadratic extension of $\mathcal{F}$, $\underline{0}$ is the zero matrix of the same size as $A$ and $r\times n$ is the size of $\left(\begin{smallmatrix}A&B\end{smallmatrix}\right)$.
@@
Note that $BB\text{${}^{\text{T}}$}=nI_{t}=CC\text{${}^{\text{T}}$}$ for some $t$. Use 00 for a zero matrix whose size is clear from the context. Now $(\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)+i\left(\begin{smallmatrix}0\\ C\end{smallmatrix}\right)z)(iC\text{${}^{\text{T}}$}+B\text{${}^{\text{T}}$}z)=0+\left(\begin{smallmatrix}0\\ I_{t}\end{smallmatrix}\right)-\left(\begin{smallmatrix}0\\ I_{t}\end{smallmatrix}\right)+0=0$ and so $H\text{${}^{\text{T}}$}(z)=iC\text{${}^{\text{T}}$}+B\text{${}^{\text{T}}$}z$ is a control matrix for the code. Hence $H(z^{-1})z=B+iC$ generates the dual of the code. It is easy then to see then that the code is dual-containing. Also a right inverse for $G(z)$ is readily written down and so the code is non-catastrophic. ∎
@@
Let $\mathcal{F}=GF(p)$ where $p\nmid n$, $A$ consist of $r$ rows of $H$ and $B$ consist of the other $(n-r)$ rows of $H$. Define $G(z)=A+iB_{1}z$ where $B_{1}=\left(\begin{smallmatrix}0_{t\times n}\\ B\end{smallmatrix}\right)$, $t=2r-n$ and $i=\sqrt{-1}$ in $\mathcal{F}$ or in a quadratic extension of $\mathcal{F}$. Then $G(z)$ generates a dual-containing convolutional code $(n,r,n-r,1,d)$.
The distance $d$ in Algorithm 4 can be estimated from Proposition 2.4 as follows: Let $A_{1}$ be the matrix of the first $(2r-n)$ of $A$ in Proposition 2.4; the distance of $\mathcal{C}$ is then $\min\{d(A_{1}),d(A)+d(\left(\begin{smallmatrix}A_{1}\\ B\end{smallmatrix}\right))\}$.
@@
The Hadamard matrix over $GF(3)$ has entries $\{1,-1\}$ which are all the non-zero entries of $GF(3)$. $GF(5)$ has the property that it contains a square root of $(-1)$ as $2=\sqrt{-1}$ in $GF(5)$. But also $GF(3)$ may be extended to $GF(3^{2})$ which has a square root of $(-1)$; the significance of $\sqrt{-1}$ is clear from the convolutional codes derived as in Propositions 2.3 and 2.4.
@@
(Proposition 3.8 in [38].) Let $\mathcal{C}$ be the code generated by $G=(I_{n},P)$. Suppose the code generated by any $s$ rows of $P$ has distance $\geq(d-s)$ and for some choice of $r$ rows the code generated by these $r$ rows has distance exactly $(d-r)$, then the distance of $\mathcal{C}$ is $d$.
@@
Modulo $p$, $HH\text{${}^{\text{T}}$}=nI_{n}=0_{n\times n}$. If $A$ is an $m\times n$ matrix and $B$ is $n\times k$, then $\operatorname{rank}(A)+\operatorname{rank}(B)-n\leq\operatorname{rank}(AB)$. Let $H=A,B=H\text{${}^{\text{T}}$}$. Then $\operatorname{rank}(H)+\operatorname{rank}(H\text{${}^{\text{T}}$})-n\leq\operatorname{rank}HH\text{${}^{\text{T}}$}=0$. But $\operatorname{rank}(H)=\operatorname{rank}(H\text{${}^{\text{T}}$})$ and so $2\operatorname{rank}(H)\leq n$ and hence $\operatorname{rank}(H)\leq\frac{n}{2}$. ∎
@@
Let $H=H(2n)$ be a Hadamard matrix of size $2n$ and $p\mid n$, $p$ a prime, $p\neq=2$. Suppose over $GF(p)$ that $H$ has rank $n$ and let $A$ be an $n\times 2n$ submatrix of rank $n$. Then $A$ generates a self-dual $[2n,n]_{p}$ code over $GF(p)$.
@@
It may be assumed that $A$ consists of the first $n$ rows of $H$ as interchanging rows of a Hadamard matrix results in a Hadamard matrix of the same rank. Thus over $GF(p)$, $HH\text{${}^{\text{T}}$}$ has the form $\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)\left(\begin{smallmatrix}A\text{${}^{\text{T}}$}&B\text{${}^{\text{T}}$}\end{smallmatrix}\right)=0$. Thus $AA\text{${}^{\text{T}}$}=0$. Now $\operatorname{rank}(A)=n$ and thus $\operatorname{rank}(A\text{${}^{\text{T}}$})=n$. Hence $A$ generates a $[2n,n]$ code $\mathcal{A}$ and $A\text{${}^{\text{T}}$}$ is a control matrix of this code. Thus the dual code of $\mathcal{A}$ is generated by $A$ and so $A$ is self-dual. ∎
@@
  • Over $GF(5)$ the code generated by $A$ is an LCD code as is the dual code is generated by $B$. Both are $[12,6,6]_{5}$ codes over $GF(5)$.
    Over $GF(5)$ define $G(z)=A+iBz$ where $i=\sqrt{-1}$; in this case $i=2$ as $12^{2}=4=-$ 1 in $GF(5)$. $G(z)$ generates a convolutional memory $1$ code which is non-catastrophic as $(A+iBz)A\text{${}^{\text{T}}$}=6I_{6}=I_{6}$ so that $A+iBz$ has a right polynomial inverse. Now $(A+iBz)(iB\text{${}^{\text{T}}$}+A\text{${}^{\text{T}}$}z)=0$ so $K\text{${}^{\text{T}}$}(z)=iB\text{${}^{\text{T}}$}+A\text{${}^{\text{T}}$}z$ is a control matrix giving that $K(z^{-1})z=A+iBz$ generates the dual code code and so the code is self-dual. The free distance of the code is the sum of the distances of the codes generated by $A$ and by $B$ which is $12$, see Proposition 2.3. Thus a self-dual convolutional $(12,6,6;1,12)_{5}$ code is obtained. From this a quantum error-correcting convolutional code is obtained with length $12$ and distance $12$ over $GF(5)=\mathbb{Z}_{5}$.
@@
  • In item 2.9 require that $01+2\alpha^{2}=$ 0 in characteristic $5$ which requires $42\alpha^{2}=-1=$ 4 which requires $\alpha^{2}=2$. Now $x^{2}-2$ is irreducible over $GF(5)$ and so extend $GF(5)$ to $GF(5^{2})$ which has an element $\alpha^{2}=2$. Then over this field $(I,\alpha H)$ generates a self-dual code. The length of the code turns out to be $8$ and thus get a $[24,12,8]$ self-dual code over $GF(5^{2})$.
    $\alpha=2$ in $GF(7)=\mathbb{Z}_{7}$ satisfies $01+12\alpha^{2}=$ 0 and so $(I,2H)$ generates a self-dual $[24,12,8]_{7}$ code.
@@
Hadamard matrices can be used to construct higher memory convolutional codes. Let $H$ be a Hadamard $12\times 12$ matrix, $A=H[1.3][1..12],B=H[4..6][1..12],C=H[7..9][1..12],D=H[10..12][1..12]$. Then $G(z)=A+Bz+Cz^{2}+Dz^{3}$ gives a $(12,3,9;3,24)$ convolutional code. The distance is easily computed as $d(A)=6=d(B)=d(C)=d(D)$ and $d(\left(\begin{smallmatrix}X\\ Y\end{smallmatrix}\right)=6$ for $X,Y$ different elements of $\{A,B,C,D\}$.
@@
With $H=H(72)$ any $36$ rows generate a $[72,36,18]_{3}$ self-dual code. With $H=H[144]$ over $GF(3)$, $72$ rows of $H$ generate a $[144,72,d]_{3}$ code.
@@
Over $GF(5)$ a $[20,10,8]_{5}$ self-dual code is obtained from $H$. But also $A=H{[1..10]}{[1..20]}$ and $B=H{[11..20]}{[1..20]}$ give $[20,10,8]_{5}$ codes over $GF(5)$.

### v2

# On codes induced from Hadamard matrices  [^1] Keywords: Hadamard, code, linear, convolutional, self-dual, dual-containing, quantum code, linear complementary AMS Classification 2020: 94B05, 94B10, 15B99

@@ 1 Introduction
$C^{\perp}$ denotes the dual of the code $\mathcal{C}$. See section 1.1 for precise definition of dual of a convolutional code. $\mathcal{C}$ is a dual-containing (DC) code if $\mathcal{C}\cap C^{\perp}=\mathcal{C}^{\perp}$; $\mathcal{C}$ is a linear complementary dual (LCD) code if $\mathcal{C}\cap C^{\perp}=0$. $\mathcal{C}$ is a self-dual code provided $\mathcal{C}^{\perp}=\mathcal{C}$ and is an important type of dual-containing code. Dual-containing codes, which include self-dual codes, can be used to construct quantum error-correcting codes (QECC) by the CSS method [13, 14, 39]; here convolutional quantum error-correcting codes are constructed in this way from Hadamard matrices.
A Hadamard matrix is an $n\times n$ matrix $H$ with entries $\pm 1$ satisfying $HH\text{${}^{\text{T}}$}=nI_{n}$. Such a matrix can only exist for $n=2$ or $n=4m$ for a positive integer $m$, [5] Theorem 18.1; see [6] for a beautifully written book on Hadamard matrices. Here these $\pm 1$ entries may be considered as elements in a general field.
The Walsh-Hadamard codes can be formed from Walsh-Hadamard matrices of size $2^{k}\times 2^{k}$ by equating the entries (1)(-1) to 0 and then forming binary codes from the $(k+1)$ linearly independent rows remaining. Using the unit-derived methods on a general Hadamard $n\times n$ matrix gives much more scope, arbitrary rates, good distances, required types but also both linear block and convolutional codes are formed. Several algorithms exist for decoding convolutional codes, the most common ones being the Viterbi algorithm and the sequential decoding algorithm.
@@
Applications are obtained by applying the Propositions and Algorithms to particular cases. Once the Hadamard matrix is formed, the Propositions and Algorithms may then be applied to produce multiple cases of required codes.
@@
    self-dual $20,10,8]_{5}$ codes;
@@
    $40,20,12]_{3}$ self-dual codes given directly in systematic form, see Proposition 2.2.
@@
  • LCD $28,16,6]_{3},[28,12,9]_{5}$ codes;
  • Convolutional self-dual $(28,14,14;1,12)_{3}$ codes over $GF(3)$;
  • DC convolutional $(28,18,10;1,8)$ codes over $GF(3)$;
  • quantum codes of length $28$, distance $8$, rate $\frac{8}{28}$ over $GF(3)$;
  • self-dual convolutional $(28,14,14;1,16)$ codes over $GF(5)$
  • DC convolutional $(28,16,12;1,14)$ codes over $GF(5)$;
  • $28,14,9]_{7}$ self-dual codes.
Generally from $H(n)$ with $p\nmid n$, self-dual $(n,\frac{n}{2},\frac{n}{2};1,d)$ convolutional codes and DC (n,r,nr;1;d)(n,r,n-r;1;d), $r>\frac{n}{2}$, convolutional codes are formed. In prototype example 2.9, it is shown how the different types of LCD, self-dual, DC, quantum, linear block and convolutional codes may be derived for a small Hadamard matrix case. Self-dual codes over $GF(p)$ can often be obtained from a Hadamard matrix of size $n$ when $p\mid n,p\neq 2$, see Proposition 2.8. For example self-dual $12k,6k,d]_{3}$ codes are produced from Paley-Hadamard matrices of size $12k$, see section 2.1.
@@ 1.1 Additional notation and background
The notation for linear block codes is standard and may be found in [1, 2, 3, 4] and many others. $GF(q)$ denotes the finite field with $q$ elements and $\mathbb{Z}_{m}$ denotes the integers modulo $m$; in particular $\mathbb{Z}_{p}=GF(p)$ for a prime $p$. A $[n,r,d]$ code denotes a linear block code of length $n$, dimension $r$, and (minimum) distance $d$; the rate is $\frac{r}{n}$. A $n,r,d]_{q}$ code denotes a linear block code of length $n$, dimension $r$ and distance $d$ over the field $GF(q)$.
Different equivalent definitions for convolutional codes are given in the literature. The notation and definitions used here follow that given in [35, 38, 36]. A rate $\frac{k}{n}$ convolutional code with parameters $(n,k,\delta)$ over a field $\mathcal{F}$ is a submodule of $\mathcal{F}[z]^{n}$ generated by a reduced basic matrix $G[z]=(g_{ij})\in\mathcal{F}[z]^{r\times n}$ of rank $r$ where $n$ is the length, $\delta=\sum_{i=1}^{r}\delta_{i}$ is the degree with $\delta_{i}=\max_{1\leq j\leq r}{\deg g_{ij}}$. Then $\mu=\max_{1\leq i\leq r}{\delta_{i}}$ is known as the memory of the code and the code is then given with parameters $(n,k,\delta;\mu)$. The parameters $(n,k,\delta;\mu,d_{f})$ are used for such a code with free (minimum) distance $d_{f}$. Further $(n,k,\delta;\mu,d_{f})_{q}$ is used to specify that the code is over the field $GF(q)$.
@@
[…]
@@ 2 Linear block and convolutional codes induced from Hadamard matrices
The Walsh-Hadamard binary linear block codes $2^{k},k,2^{k-1}]_{2}$ and $2^{k},k+1,2^{k-1}]_{2}$ have very small rate but have found use probably on account of the distances and decoding methods available. Codes from general Hadamard matrices as now described allow much more scope with much better rates, good distances, required types and both linear block and convolutional codes may be formed.
@@
Let $H$ be a Hadamard matrix of size $n$ so that $HH\text{${}^{\text{T}}$}=nI_{n}$ and $n\neq 0$ in a field $\mathcal{F}$. Suppose $H$ has the form $H=\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)$ implying $\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)\left(\begin{smallmatrix}A\text{${}^{\text{T}}$}&B\text{${}^{\text{T}}$}\end{smallmatrix}\right)=nI_{n}$ where $A$ has size $r\times n$ and $B$ has size $(n-r)\times n$ with r>(nr)r>(n-r). Let t=(2rn)t=(2r-n) and define $B_{1}=\bigl(\begin{smallmatrix}0_{t\times n}\\ B\end{smallmatrix}\bigr)$. Then $G(z)=A+iB_{1}z$ generates a convolutional dual-containing (n,r,nr;1,d)(n,r,n-r;1,d) code $\mathcal{C}$ where $i=\sqrt{-1}$ in $\mathcal{F}$ or in a quadratic extension of $\mathcal{F}$.
@@
Define $0_{t}=0_{t\times n}$. Thus $B_{1}=\left(\begin{smallmatrix}0_{t}\\ B\end{smallmatrix}\right)$ is an $r\times n$ matrix. Now $A\text{${}^{\text{T}}$}$ is an $n\times r$ matrix and thus has the form $A\text{${}^{\text{T}}$}=(X,C_{1})$ where $C_{1}$ has size $n\times(n-r)$ and $X$ has size $n\times(2r-n)$. As $AA\text{${}^{\text{T}}$}=nI_{r}$ then $AC_{1}=n\begin{pmatrix}0_{(2r-n)\times(n-r)}\\ I_{(n-r)\times(n-r)}\end{pmatrix}$ and also $B_{1}B=n\begin{pmatrix}0_{(2r-n)\times(n-r)}\\ I_{(n-r)\times(n-r)}\end{pmatrix}$. Now $A=\left(\begin{smallmatrix}X\text{${}^{\text{T}}$}\\ C_{1}\text{${}^{\text{T}}$}\end{smallmatrix}\right),B_{1}=\left(\begin{smallmatrix}0_{t}\\ D\text{${}^{\text{T}}$}\end{smallmatrix}\right)$. Then $(A+iB_{1}z)(iB\text{${}^{\text{T}}$}+C_{1}z)=0$ so $H\text{${}^{\text{T}}$}(z)=(iB\text{${}^{\text{T}}$}+C_{1}z)$ is a control matrix and $H(z^{-1})z=C_{1}\text{${}^{\text{T}}$}+iB$ generates the dual of the code. The code generated by $C_{1}\text{${}^{\text{T}}$}+iB$ is easily seen to be contained in the code generated by $(A+iB_{1})$ and so the code generated by $(A+iB_{1})$ is dual-containing. That the code is non-catastrophic follows in a similar manner to the proof in Proposition 2.3.
@@
Then $G(z)=\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)+i\left(\begin{smallmatrix}\underline{0}\\ C\end{smallmatrix}\right)$ defines a dual-containing convolutional (n,r,nr;1,d)(n,r,n-r;1,d) code where $i=\sqrt{-1}$ in $\mathcal{F}$ or in a quadratic extension of $\mathcal{F}$, $\underline{0}$ is the zero matrix of the same size as $A$ and $r\times n$ is the size of $\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)$.
@@
Note that $BB\text{${}^{\text{T}}$}=nI_{t}=CC\text{${}^{\text{T}}$}$ for some $t$. Use 0 for a zero matrix whose size is clear from the context. Now $(\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)+i\left(\begin{smallmatrix}0\\ C\end{smallmatrix}\right)z)(iC\text{${}^{\text{T}}$}+B\text{${}^{\text{T}}$}z)=0+\left(\begin{smallmatrix}0\\ I_{t}\end{smallmatrix}\right)-\left(\begin{smallmatrix}0\\ I_{t}\end{smallmatrix}\right)+0=0$ and so $H\text{${}^{\text{T}}$}(z)=iC\text{${}^{\text{T}}$}+B\text{${}^{\text{T}}$}z$ is a control matrix for the code. Hence $H(z^{-1})z=B+iC$ generates the dual of the code. It is easy to see that the code is dual-containing. Also a right inverse for $G(z)$ is readily written down and so the code is non-catastrophic. ∎
@@
Let $\mathcal{F}=GF(p)$ where $p\nmid n$, $A$ consist of $r$ rows of $H$ and $B$ consist of the other (nr)(n-r) rows of $H$. Define $G(z)=A+iB_{1}z$ where $B_{1}=\left(\begin{smallmatrix}0_{t\times n}\\ B\end{smallmatrix}\right)$, t=2rnt=2r-n and $i=\sqrt{-1}$ in $\mathcal{F}$ or in a quadratic extension of $\mathcal{F}$. Then $G(z)$ generates a dual-containing convolutional code (n,r,nr,1,d)(n,r,n-r,1,d).
The distance $d$ in Algorithm 4 can be estimated from Proposition 2.4 as follows: Let $A_{1}$ be the matrix of the first (2rn)(2r-n) of $A$ in Proposition 2.4; the distance of $\mathcal{C}$ is then $\min\{d(A_{1}),d(A)+d(\left(\begin{smallmatrix}A_{1}\\ B\end{smallmatrix}\right))\}$.
@@
The Hadamard matrix over $GF(3)$ has entries $\{1,-1\}$ which are all the non-zero entries of $GF(3)$. $GF(5)$ has the property that it contains a square root of (1)(-1) as $2=\sqrt{-1}$ in $GF(5)$. But also $GF(3)$ may be extended to $GF(3^{2})$ which has a square root of (1)(-1); the significance of $\sqrt{-1}$ is clear from the convolutional codes derived as in Propositions 2.3 and 2.4.
@@
(Proposition 3.8 in [27].) Let $\mathcal{C}$ be the code generated by $G=(I_{n},P)$. Suppose the code generated by any $s$ rows of $P$ has distance $\geq(d-s)$ and for some choice of $r$ rows the code generated by these $r$ rows has distance exactly (dr)(d-r), then the distance of $\mathcal{C}$ is $d$.
@@
[…]
@@
Let $H=H(2n)$ be a Hadamard matrix of size $2n$ and $\mid n, pp$ a prime, $p\neq 2$. Suppose over $GF(p)$ that $H$ has rank $n$ and let $A$ be an $n\times 2n$ submatrix of rank $n$. Then $A$ generates a self-dual $2n,n]_{p}$ code over $GF(p)$.
@@
It may be assumed that $A$ consists of the first $n$ rows of $H$ as interchanging rows of a Hadamard matrix results in a Hadamard matrix of the same rank. Thus over $GF(p)$, $HH\text{${}^{\text{T}}$}$ has the form $\left(\begin{smallmatrix}A\\ B\end{smallmatrix}\right)\left(\begin{smallmatrix}A\text{${}^{\text{T}}$}&B\text{${}^{\text{T}}$}\end{smallmatrix}\right)=0$. Thus $AA\text{${}^{\text{T}}$}=0$. Now $\operatorname{rank}(A)=n$ and thus $\operatorname{rank}(A\text{${}^{\text{T}}$})=n$. Hence $A$ generates a $[2n,n]$ code $\mathcal{A}$ and $A$ is a control matrix of this code. Thus the dual code of $\mathcal{A}$ is generated by $A$ and so $\cal{A}$ is self-dual. ∎
@@
  • Over $GF(5)$ the code generated by $A$ is an LCD code as is the dual code is generated by $B$. Both are $12,6,6]_{5}$ codes over $GF(5)$.
    Over $GF(5)$ define $G(z)=A+iBz$ where $i=\sqrt{-1}$; in this case $i=2$ as $2^{2}=4=-1$ in $GF(5)$. $G(z)$ generates a convolutional memory $1$ code which is non-catastrophic as $(A+iBz)A\text{${}^{\text{T}}$}=6I_{6}=I_{6}$ so that $A+iBz$ has a right polynomial inverse. Now $(A+iBz)(iB\text{${}^{\text{T}}$}+A\text{${}^{\text{T}}$}z)=0$ so $K\text{${}^{\text{T}}$}(z)=iB\text{${}^{\text{T}}$}+A\text{${}^{\text{T}}$}z$ is a control matrix giving that $K(z^{-1})z=A+iBz$ generates the dual code code and so the code is self-dual. The free distance of the code is the sum of the distances of the codes generated by $A$ and by $B$ which is $12$, see Proposition 2.3. Thus a self-dual convolutional $(12,6,6;1,12)_{5}$ code is obtained. From this a quantum error-correcting convolutional code is obtained with length $12$ and distance $12$ over $GF(5)=\mathbb{Z}_{5}$.
@@
  • In item 2.9 require that $1+2\alpha^{2}=0$ in characteristic $5$ which requires $2\alpha^{2}=-1=4$ which requires $\alpha^{2}=2$. Now $x^{2}-2$ is irreducible over $GF(5)$ and so extend $GF(5)$ to $GF(5^{2})$ which has an element $\alpha^{2}=2$. Then over this field $(I,\alpha H)$ generates a self-dual code. The length of the code turns out to be $8$ and thus get a $[24,12,8]$ self-dual code over $GF(5^{2})$.
    In $GF(7)=\mathbb{Z}_{7}$, $\alpha=2$ satisfies $1+12\alpha^{2}=0$ and so $(I,2H)$ generates a self-dual $24,12,8]_{7}$ code.
@@
Hadamard matrices can be used to construct higher memory convolutional codes. Let $H$ be a Hadamard $12\times 12$ matrix, A=H[1.3][1..12],B=H[4..6][1..12],C=H[7..9][1..12],D=H[10..12][1..12]A=H[1.3][1..12],B=H[4..6][1..12],C=H[7..9][1..12],D=H[10..12][1..12]. Then $G(z)=A+Bz+Cz^{2}+Dz^{3}$ gives a $(12,3,9;3,24)$ convolutional code. The distance is easily computed as $d(A)=6=d(B)=d(C)=d(D)$ and $d(\left(\begin{smallmatrix}X\\ Y\end{smallmatrix}\right)=6$ for $X,Y$ different elements of $\{A,B,C,D\}$.
@@
With $H=H(72)$ any $36$ rows generate a $72,36,18]_{3}$ self-dual code. With $H=H[144]$ over $GF(3)$, $72$ rows of $H$ generate a $144,72,d]_{3}$ code.
@@
Over $GF(5)$ a $20,10,8]_{5}$ self-dual code is obtained from $H$. But also A=H[1..10][1..20]A=H{[1..10]}{[1..20]} and B=H[11..20][1..20]B=H{[11..20]}{[1..20]} give $20,10,8]_{5}$ codes over $GF(5)$.